OpenStudy (anonymous):

what is the range of y=2x-2?

OpenStudy (mathstudent55):

Do you know what the range of a function is?

OpenStudy (anonymous):

I know that the domain is all real numbers and I graphed it out but I am confused on the range I believe that the range is \[1\le x \le -2\] but im not sure

OpenStudy (anonymous):

or is it \[1\le x \le 1\]

OpenStudy (mathstudent55):

You still haven't answered my question. Do you know what the range of a function is? I don't mean the range of this function. In general, what is the range of a function?

OpenStudy (anonymous):

oh it is the y value. oops those were supposed to be y's in those equations

OpenStudy (mathstudent55):

The domain of a function is the set of values that can be used for x. The range of a function is the set of values that can be used for y. Ok so far?

OpenStudy (anonymous):

Yes that is what I meant when I said that range is the y values.

OpenStudy (mathstudent55):

Ok, so you understand the concept. That's good. For the function y = 2x - 2, is there any number you can't use for x?

OpenStudy (mathstudent55):

Usually, there are restrictions on x values when there is a square root with x inside or x in a denominator. Here, there are no restrictions. x can be any real number. That means the domain is all real numbers.

OpenStudy (anonymous):

Yes I understand that part. Now how do I find the range?

OpenStudy (mathstudent55):

Ok, now that you understand that the domain is every real number, think about the range. The range is all the y values. If you substitute x by every real number, then y will also be every real number.

OpenStudy (mathstudent55):

In fact, if you graph this equation, you'll see that it is a straight line that never ends in both directions. Both the x values and y values go one forever. That means both the domain and the range are all real numbers.

OpenStudy (anonymous):

Ok so the range is also all real numbers? That makes sense. I originally thought that but for some reason I thought that the range could not be all real numbers.

OpenStudy (mathstudent55):

As you can see in the graph, as the line goes more and more up to the right, the y-coordinates will get higher and higer for ever. As you go more and more down to the left on the line, y-coordinates go lower and lower also forever. The range is all real numbers.